# BlackboxBench: A Comprehensive Benchmark of Black-box Adversarial Attacks

### v1

@@ Abstract
Adversarial examples are well-known tools to evaluate the vulnerability of deep neural networks (DNNs). Although lots of adversarial attack algorithms have been developed, it is still challenging in the practical scenario that the model’s parameters and architectures are inaccessible to the attacker/evaluator, $i.e.$, black-box adversarial attacks. Due to the practical importance, there has been rapid progress from recent algorithms, reflected by the quick increase in attack success rate and the quick decrease in query numbers to the target model. However, there is a lack of thorough evaluations and comparisons among these algorithms, causing difficulties of tracking the real progress, analyzing advantages and disadvantages of different technical routes, as well as designing future development roadmap of this field. Thus, in this work, we aim at building a comprehensive benchmark of black-box adversarial attacks, called BlackboxBench. It mainly provides: 1) a unified, extensible and modular-based codebase, implementing 25 query-based attack algorithms and 30 transfer-based attack algorithms; 2) comprehensive evaluations: we evaluate the implemented algorithms against several mainstreaming model architectures on 2 widely used datasets (CIFAR-10 and a subset of ImageNet), leading to 14,106 evaluations in total; 3) thorough analysis and new insights, as well analytical tools. The website and source codes of BlackboxBench are available at https://blackboxbench.github.io/ and https://github.com/SCLBD/BlackboxBench/, respectively.
@@ I Introduction
Deep neural networks (DNNs) have been densely applied to safety-critical applications such as autonomous driving and face recognition due to their remarkable success. Worryingly, adversarial examples, which add human-imperceptible noises to benign inputs, capable of deceiving well-trained models, pose a significant threat to DNNs. In real-world scenarios, attackers typically have limited access to information about the target classifier, making them black-box adversaries.
Machine learning (ML) security has witnessed a rapid evolution of black-box adversarial attacks. Numerous researchers strive to enhance the state-of-the-art (SOTA) attack performance against vulnerable DNNs [132, 133, 134]. These endeavors are also inherently double as robustness assessments to develop stronger defenses and more resilient models [135, 136, 137, 138]. They are like the spear and the shield, mutually reinforcing each other. However, the proliferation of diverse works, varying in experimental settings and code styles, perplexes researchers regarding the actual utiluty and security efficacy of attacks and defenses. As a result, there is a growing demand for building a comprehensive platform for black-box adversarial attacks to facilitate fair comparisons within the community. Moreover, the identification of factors contributing to adversarial vulnerabilities remains unclear, perpetuating a vicious cycle in the study of neural network robustness. Such a platform not only addresses these challenges but also empowers a more profound exploration of this topic through an analysis of its intrinsic properties.
Contributions We build a comprehensive benchmark of black-box adversarial attacks, called BlackboxBench. It provides the following key contributions:
  • Modular&unified codebase BlackboxBench takes a modular and unified implementation. As the core part of BlackboxBench, $i.e.$, the attack module, is implemented based on a unified pipeline composed of several functional blocks (see Fig. 4), BlackboxBench could be easily extended to cover more latest methods, and could also facilitate the involvement of new contributors. To date, we have included 25 query-based attacks and 30 transfer-based attacks, the largest collection of black-box adversarial attacks to our best knowledge.
  • Comprehensive evaluations Using BlackboxBench, we conduct evaluations of 25 query-based attacks and 30 transfer-based attacks on comprehensive attack scenarios, leading to 14,106 evaluations in total. These standardized evaluations establish a leaderboard to identify the most effective methods and track actual progress of this field.
  • Thorough analysis The comprehensive evaluations enable us to delve deeper into black-box adversarial attacks. We further present 13 types of analysis to reveal the mechanisms of black-box attacks and the effects of various factors to attack performance, with the assistance of ten analytical tools. The insights from our evaluations, as well as the provided analytical tools, could not only broaden and deepen our understanding of the inherent characteristics of black-box attacks, but also facilitate researchers to develop more advanced attack and defense methods.
@@
The remaining parts of this manuscript are organized as follows. Section 2 introduces the related benchmarks and the uniqueness of BlackboxBench. Section 3 provides the preliminary, introduction to implemented algorithms, an overview of BlackboxBench architecture and library modules. Section 4 provides the evaluation results, systematic analyses and our new insights.
@@ II Related Work
ML models have demonstrated vulnerability to imperceptive deliberate perturbations. Some researchers identify these vulnerabilities by designing new attacks, others pay ongoing efforts toward new defenses against such perturbations [140]. These perturbations encompass various types (e.g., patch-based, norm-bounded) to fool models in different ML tasks (e.g., computer vision, natural language understanding). Within this research community, many benchmarks have emerged for standardized evaluations. GRB [125] evaluates the adversarial robustness of graph machine learning. SmartBox [123] benchmarks the performance of adversarial attack detection and mitigation algorithms against face recognition. REAP [124] is a benchmark for security against patch attacks. AdvGLUE [126] evaluates the vulnerabilities of modern large-scale language models. CARBEN [127] and MultiRobustBench [128] are unified frameworks for considering multiple attacks against ML models.
@@
BlackboxBench is designed for vision tasks, currently the most extensively studied ML setting. Moreover, unlike some excellent benchmarks specialized in defenses (e.g., RobustBench, [42], RobustML), our aim is to track the real progress in black-box adversarial attacks. Various libraries including CleverHans[82], FoolBox[81], ART[85], DEEPSEC [130] AdverTorch[83], SecML[86], RealSafe [131], AdvBox[84], DeepRobust[87], and TransferAttackEval [129] have all provided evaluation for popular adversarial attacks. Nevertheless, BlackboxBench sets itself apart from existing benchmarks from five key perspectives, as detailed in Tab. I.
  • Programming characteristic Most libraries employ the object-oriented programming (OOP) language. However, we notice that all the attack algorithms within query-based category or transfer-based category share highly similar procedures. In essence, all the methods can be implemented through specific modules (see Fig. 1). Particularly, their modules for generating perturbations ($i.e.$, attack module) can be consolidated into a unified pipeline, comprising multiple functional blocks (see Fig. 4). By building a modular and unified procedural programming structure, BlackboxBench can easily switch from one method to another by just modifying a single functional block, without redefining the whole optimization process like OOP-style benchmarks. Note that while DEEPSEC and SecML also feature a modular architecture, they lack an unified attack module compared to BlackboxBench.
  • Code framework BlackboxBench is implemented in PyTorch, a widely used framework within the deep learning research community.
@@
  • Algorithms coverage The number and published years of white-box (WA) and black-box (BA) adversarial attacks  [^1] White-box attacks can tackle black-box tasks owing to the transferability of adversarial examples. Without specifications in original papers, we categorize white-box attacks evaluated within a black-box setting as black-box attacks. that each library covers is listed in the fifth and sixth column of Tab. I, respectively. BlackboxBench stands out by supporting 55 black-box attack methods from 2017 to 2023, diverging from other benchmarks that mainly encompass white-box or early black-box techniques.
@@
  • Leaderboard BlackboxBench is the first to provide an up-to-date leaderboard to track the progress in black-box adversarial attacks.
@@
  • Analysis & tools Besides analyzing quantitative results like DEEPSEC, RealSafe and TransferAttackEval, BlackboxBench also offers several analysis tools to elucidate the underlying mechanisms of generating adversarial samples in black-box scenarios, striving to provide fresh perspectives for future research in this domain. Note that although SecML provide tools, only partial tools are involved in the analysis presented in its paper.
TABLE I: Comparison between BlackboxBench and other benchmarks.
[…]
CleverHans[82]
A library provides standardized implementations of attacks and adversarial training.
All attacks are independent of each other.
JAX, PyTorch, Tensorflow
[…]
FoolBox[81]
A collection of gradient-based and decision-based adversarial attacks.
Object-oriented programming.
PyTorch, Tensorflow, JAX
[…]
ART[85]
ART provides tools to defend and evaluate ML models against the Evasion, Poisoning, Extraction, and Inference attacks.
Object-oriented programming.
All popular ML frameworks
[…]
DEEPSEC[130]
DEEPSEC incorporates attacks, defenses, and multiple utility metrics.
All modules are independent of each other, but multiple modules can also work together.
PyTorch 0 / 16 2014-2018 \
[…]
AdverTorch[83]
AdverTorch contains modules for attacks and defenses, also scripts for robust training
[…]
SecML[86]
SecML implements: (i) evasion attacks and poisoning attacks; and (ii) explainable methods to understand attacks.
SecML has a modular architecture oriented to code reuse. Each module owns an interface.
PyTorch
Attacks in CleverHans and Foolbox
[…]
RealSafe[131]
RealSafe benchmarks the adversarial robustness of image classification and object detection models, and introduces robust training.
Object-oriented programming. PyTorch 5 / 10 2015-2019 \
[…]
AdvBox[84]
AdvBox supports black box attacks on Machine-Learning-as-a-service, as well as more attack scenarios, such as Face Recognition Attack.
Object-oriented programming.
PaddlePaddle, Keras, PyTorch, TensorFlow, Caffe2, MxNet
[…]
DeepRobust[87]
DeepRobust contains attacks and defenses in image domain and graph domain.
[…]
TransferAttackEval[129]
The first large-scale evaluation of transferable adversarial examples.
All attacks are independent of each other.
PyTorch 0 / 23 2018-2022 \
[…]
BlackboxBench
A comprehensive benchmark of black-box adversarial attacks, including query-based attacks and transfer-based attacks.
Modular, unified procedural programming.
PyTorch 0 / 55 2017-2023 BlackboxBench LeaderBoard
[…]
@@ III-A Preliminary
BlackboxBench focuses on black-box adversarial attacks against deep learning models. This section seeks to provide the novice reader with a complete grasp of black-box adversarial attacks.
Given the victim neural network $f:\mathbb{R}^{n}\rightarrow\mathcal{P}(\mathcal{Y})$, parameterized by $\theta$, where $\mathcal{P}(\mathcal{Y})$ is the $K$-dimensional probability simplex. For a benign example $\boldsymbol{x}$ whose ground truth label is $y$, adversarial attacks aim at misleading, with imperceptible perturbation $\boldsymbol{\delta}_{\boldsymbol{x}}$, the prediction of $\boldsymbol{x}$ from $f$, which lead to the adversarial example $\boldsymbol{x}^{*}=\boldsymbol{x}+\boldsymbol{\delta}_{\boldsymbol{x}}$. Formally, $\boldsymbol{x}^{*}$ satisfies two properties:
  • To make sure $\boldsymbol{x^{*}}$ is perceptually similar to $\boldsymbol{x}$, $\boldsymbol{x^{*}}$ is generated within a $\ell_{p}$ norm ball certered at $\boldsymbol{x}$ with radius $\epsilon$, $i.e.$, $\left\|\boldsymbol{x^{*}}-\boldsymbol{x}\right\|_{p}\leq\epsilon$. $\ell_{\infty}$ and $\ell_{2}$ are the most commonly used norms. The perturbation norm budget $\epsilon$ is small enough so as to remain undetectability.
@@
Depending on the full access to the $f$ or not, adversarial attacks can be divided into white-box attacks and black-box attacks. The former is deemed easier but less realistic because attackers could access parameters of the victim models. The latter is more practical and challenging because no information is known about the victim, and is the focus of BlackboxBench. As background, we introduce different types of black-box attacks: query-based attacks, which query the target model repeatedly and can be further categorized into score-based and decision-based according to the model feedback, and transfer-based attacks, which work offline without interacting with the target model222[122] categorize adversarial attacks as white-box, black-box (specifically referring to query-based attacks) and transfer-based attacks. Here we incorporate transfer-based attacks into black-box category to emphasize the importance of adversarial transferability in enhancing black-box attack power..
@@ III-A1 Decision-based query attacks
Decision-based query attacks primarily work with only the final decision of a model (e.g., the classification output) without access to the model’s gradients, logits, or any other internal states. Their general formulation is:
$\min_{\boldsymbol{x^{*}}}\mathcal{D}\left(\boldsymbol{x},\boldsymbol{x^{*}}\right)+\xi\left(\Omega\left(f\left(\boldsymbol{x^{*}}\right),y\right)=1\right).$ (1)
The perturbation aims to be as imperceptible as possible (corresponding to minimizing the distance, represented by $\mathcal{D}$, between the initial input $\boldsymbol{x}$ and its perturbed counterpart $\boldsymbol{x^{*}}$) while keep misleading the model into incorrect classifications. The term $\xi$ introduces this hard constraint to decision-based optimization process. $\xi(u)=0$ if $u$ is ture. If not, $\xi(u)=\infty$. Within its scope, $\Omega$ symbolizes the adversarial criteria, which essentially forces the perturbed input to meet the attack goal. Depending on the type of attack, this criterion varies:
@@
where $\mathbb{I}$ is the indicator function. In essence, attackers are navigating the adversarial space, defined by the constraints of the model and the criteria. Their goal is to find a point that is both adversarial (misclassified) and as close as possible to the original input. The graphical illustration of untargeted decision-based attacks is shown in Fig. 2(a).

### v2

@@ Abstract
Adversarial examples are well-known tools to evaluate the vulnerability of deep neural networks (DNNs). Although lots of adversarial attack algorithms have been developed, it’s still challenging in the practical scenario that the model’s parameters and architectures are inaccessible to the attacker/evaluator, $i.e.$, black-box adversarial attacks. Due to the practical importance, there has been rapid progress from recent algorithms, reflected by the quick increase in attack success rate and quick decrease in query numbers to the target model. However, there lacks thorough evaluations and comparisons among these algorithms, causing difficulties in tracking the real progress, analyzing advantages and disadvantages of different technical routes, as well as designing future development roadmap of this field. Thus, we aim at building a comprehensive benchmark of black-box adversarial attacks, called BlackboxBench. It mainly provides: 1) a unified, extensible and modular-based codebase, implementing 29 query-based attack algorithms and 30 transfer-based attack algorithms; 2) comprehensive evaluations: we evaluate the implemented algorithms against several mainstreaming model architectures on 2 widely used datasets (CIFAR-10 and a subset of ImageNet), leading to 14,950 evaluations1 in total; 3) thorough analysis and new insights, as well analytical tools. The website and source codes of BlackboxBench are available at https://blackboxbenchmark.github.io/ and https://github.com/SCLBD/BlackboxBench/, respectively.
@@ I Introduction
Deep neural networks (DNNs) have been densely applied to safety-critical applications such as autonomous driving and face recognition due to their remarkable success. Worryingly, adversarial examples, which add human-imperceptible noises to benign inputs, capable of deceiving well-trained models, pose a significant threat to DNNs. In real-world situations, attackers typically have limited access to information about the target model. This makes the study of adversarial attacks with in a black-box scenario crucial for understanding the robustness of models against more realistic attack conditions.
The field of machine learning security has witnessed a rapid evolution of black-box adversarial attacks. Due to the lack of access to target model’s complete information, attackers must rely on either external queries or surrogate models, corresponding to query-based and transfer-based attacks, respectively. As a result, significant efforts have been made in these two areas to enhance the state-of-the-art (SOTA) attack performance [1, 2, 3, 4]. These endeavors are also inherently double as robustness assessments to develop stronger defenses and more resilient models [5, 6, 7, 8]. However, the proliferation of diverse attack methods, evaluated in different model architectures and datasets, compared to limited previous methods, perplexes researchers regarding the actual threat posed by them, as well as the actual progress in this field. Meanwhile, variations in code styles and lack of a unified attack pipeline make it hard to use them to examine the actual protection provided by robust models or defense mechanisms. As a result, there is a growing demand for building a comprehensive platform for black-box adversarial attacks to facilitate fair comparisons within the community.
To this end, we build a comprehensive benchmark of black-box adversarial attacks, called BlackboxBench. To date, we have included 29 query-based attacks and 30 transfer-based attacks, the largest collection of black-box adversarial attacks to our best knowledge. BlackboxBench adopts a modular and unified implementation. As the core module of BlackboxBench, $i.e.$, the attack module, is implemented based on a unified attack pipeline composed of several functional blocks (see Fig. 4), BlackboxBench allows fair and standard evaluation, illuminates promising technical routes, facilitates the involvement of new contributors, and is user-friendly. Using BlackboxBench, we conduct evaluations of all these query-based attacks and transfer-based attacks on comprehensive attack scenarios—varying attack settings, target models, and datasets—resulting in 14,950 evaluations111 One evaluation is conducted by generating adversarial examples on a dataset using a specific attack method against a particular target model under a defined attack setting. in total. These standardized evaluations establish a leaderboard to identify the most effective methods and track actual progress in this field. Our results indicate that black-box adversarial attacks have evolved significantly over the years, with CISA emerging as the SOTA method for decision-based attacks, MCG for score-based attacks, and SIA and Bayesian Attack for untargeted and targeted transfer-based attacks, respectively. Additional observations addressing benchmark-level questions (see the first four entries in Tab. I) are briefly recapped at the end of each subsection in Sec. IV-B.
Beyond benchmarking, we further leverage BlackboxBench to foster a deeper understanding of black-box adversarial vulnerability. In addition to the evaluations above, we extend our experimental scope to systematically assess the influence of critical factors-including the data used for attacking, the architecture of surrogate and target models, constraints on attack budget, deployed defenses, and attack strategies—on the black-box adversarial vulnerability of models. Unlike the benchmark evaluations, which focus on tracking progress, these additional analyses aim to uncover underlying factors behind black-box adversarial vulnerability. These analyses are supported by 10 analytical tools (see Tab. III in the Supplemental Material for details on tools and their functionalities), and are organized around 11 key analytical questions (see the latter entries in Tab. I) from the perspectives of data, model architecture, budget, defense, and attack procedure, providing several insightful takeaways, including (but not limited to) the following:
  • Data We find that samples far from the model’s decision boundary would take higher query costs for successful attacks across various query-based methods, as they make it harder for attackers to efficiently find the attack direction.
  • Model architecture We show that complex models like ConvNeXt-T and ViT-B/16, with smoother decision boundaries, exhibit stronger robustness to adversarial queries, while simpler models like ResNet-50 and VGG-19, with more fragmented decision boundaries, are more vulnerable. We also propose a quantitative metric to measure the transferability of adversarial examples between models.
  • Attack budget We find that surrogate-assisted score-based methods ($e.g.$, MCG, PRGF, BASES, $\mathcal{CG}$-attack) are more efficient in query and norm budgets. Transfer-based methods aiming for semantically meaningful perturbations ($e.g.$, FIA, NAA, DRA, IAA) generally require higher norm budgets to maintain their strong performance.
  • Defense We observe that adversarial examples transfer more effectively between models with the same training strategy, due to their shared attention in the frequency domain.
  • Attack procedure We visualize successful adversarial attacks would manage to shift the model’s attention from discriminative areas to trivial areas, thus resulting in misclassification.
Further insights into other key questions are presented at the beginning of each analytical section ($i.e.$, Sec. IV-C, IV-D and IV-E, Supplementary C-C and C-D).
@@
This paper is organized as follows. Sec. II introduces the related benchmarks and highlights the uniqueness of BlackboxBench. Sec. III provides the preliminary and an overview of BlackboxBench architecture and library modules. Sec. IV presents the evaluation results, systematic analyses and our new insights, followed by conclusions in Sec. V.
@@ II Related Work
ML models have demonstrated vulnerability to imperceptive deliberate perturbations. Some researchers demonstrate these vulnerabilities by designing new attacks, while others pay ongoing efforts toward new defenses to enhance robustness [9]. Within this research community, many benchmarks have emerged for standardized evaluations. The evaluated perturbations are of various types (e.g., patch-based[10], norm-bounded[11, 12]) to fool models in different ML tasks (e.g., computer vision[13], natural language understanding[14], [15]).
@@
BlackboxBench is designed for norm-bounded noise perturbed vision tasks, currently the most extensively studied ML setting. Moreover, unlike some excellent benchmarks specialized in defenses (e.g., RobustBench [16], RobustML), our aim is to track the real progress in black-box adversarial attacks. Various libraries including CleverHans[17], FoolBox[18], ART[19], DEEPSEC [20] AdverTorch[21], SecML[22], RealSafe [23], AdvBox[24], DeepRobust[25], TransferAttackEval [26], Torchattacks[27] and TA-Bench[28] have all provided evaluation for popular adversarial attacks. Nevertheless, BlackboxBench sets itself apart from existing benchmarks from four key perspectives, as detailed in Tab. II of the Supplemental Material.
  • Programming characteristic Most libraries employ object-oriented programming (OOP) to define attacks as separate classes. In contrast, BlackboxBench features a modular codebase with an attack module that integrates various attacks into a unified attack pipeline composed of several functional blocks. This procedural structure allows easy switching between attacks or integrating multiple attacks into a more powerful one by modifying or combining functional blocks within the unified attack pipeline, without the need to redefine the entire optimization process as required in OOP-style benchmarks. Note that while DEEPSEC[20] and SecML[22] also feature modular architectures, they lack a unified attack module compared to BlackboxBench.
@@
  • Algorithm coverage Taking a look at the number and published years of white-box (WA) and black-box (BA) adversarial attacks [^2] White-box attacks can tackle black-box tasks owing to the transferability of adversarial examples. Without specifications in original papers, we categorize white-box attacks evaluated within a black-box setting as black-box attacks. that each library covers, BlackboxBench stands out by supporting 59 black-box attack methods from 2017 to 2023, diverging from other benchmarks that mainly encompass white-box or early black-box techniques.
@@
  • Leaderboard BlackboxBench is the first platform that furnishes a leaderboard for monitoring advancements in black-box adversarial attacks.
@@
  • Tools & Analyses Besides only analyzing quantitative results like DEEPSEC, RealSafe, TransferAttackEval and TA-Bench, BlackboxBench also offers analytical tools to facilitate in-depth analyses of underlying factors behind black-box adversarial vulnerability, aiming at provide fresh perspectives for future research. Note that although SecML provides tools, only one analysis has been conducted.
@@ III-A Preliminary
BlackboxBench focuses on black-box adversarial attacks against deep learning models. This section seeks to provide novice readers with a complete grasp of this domain.
Given target neural network $f:\mathcal{X}\rightarrow\mathcal{P}(\mathcal{Y})$, parameterized by $\theta$, where $\mathcal{X}\subset\mathbb{R}^{d}$, and $\mathcal{P}(\mathcal{Y})$ is the $K$-dimensional probability simplex over $\mathcal{Y}$. For benign example $\boldsymbol{x}\in\mathcal{X}$ whose true label is $y\in\mathcal{Y}$, adversarial attacks aim at misleading, with imperceptible perturbation $\boldsymbol{\delta}_{\boldsymbol{x}}$, the prediction of $\boldsymbol{x}$ from $f$, which lead to the adversarial example $\boldsymbol{x}^{*}=\boldsymbol{x}+\boldsymbol{\delta}_{\boldsymbol{x}}$. Formally, $\boldsymbol{x}^{*}$ satisfies two properties:
  • To make sure $\boldsymbol{x^{*}}$ is perceptually similar to $\boldsymbol{x}$, $\boldsymbol{x^{*}}$ is generated within a $\ell_{p}$ norm ball centered at $\boldsymbol{x}$ with radius $\epsilon$, $i.e.$, $\left\|\boldsymbol{x^{*}}-\boldsymbol{x}\right\|_{p}\leq\epsilon$. $\ell_{\infty}$ and $\ell_{2}$ are the most commonly used norms. The perturbation norm budget $\epsilon$ is small enough so as to remain undetectability.
@@
Depending on the full access to the $f$ or not, adversarial attacks can be divided into white-box and black-box. The former is deemed easier but less realistic because attackers could access parameters of the target models. The latter is more practical and challenging because no information is known about the target model, and is the focus of BlackboxBench. As background, we introduce different types of black-box attacks: query-based attacks, which query the target model repeatedly and can be categorized into score-based and decision-based according to the model feedback, and transfer-based attacks, which work offline without interacting with the target model [^3] In the latest survey [29], adversarial attacks are categorized as white-box, black-box (specifically referring to query-based attacks) and transfer-based attacks. Here we incorporate transfer-based attacks into black-box category to emphasize the importance of adversarial transferability in enhancing black-box attack ability.. Their graphical illustrations are shown in Fig. 2.
@@ III-A1 Decision-based query attacks
Decision-based query attacks primarily work with only the final decision of a model (e.g., the classification output) without access to the model’s gradients, logits, or any other internal states. The general formulation is:
$\min_{\boldsymbol{x^{*}}}\mathcal{D}\left(\boldsymbol{x},\boldsymbol{x^{*}}\right)\text{ s.t. }\Omega\left(f\left(\boldsymbol{x^{*}}\right),y\right)=1.$ (1)
The perturbation aims to be as imperceptible as possible (corresponding to minimizing the distance, represented by $\mathcal{D}$, between the initial input $\boldsymbol{x}$ and its perturbed counterpart $\boldsymbol{x^{*}}$), while keeping misleading the model into incorrect classifications (corresponding to the hard constraint introduced by $\Omega$). $\Omega$ forces the perturbed input to meet the attack goal, which takes 1 if the constraint is satisfied, otherwise 0. According to the type of attack, the $\Omega$ is specified as follows:
@@
$\mathbb{I}(\cdot)$ is the indicator function, and $\mathbb{I}(c)=1$ if $c$ is true and $\mathbb{I}(c)=0$ otherwise. In essence, attackers are navigating the adversarial space defined by the hard constraints. Their goal is to find a point that is both adversarial (misclassified) and as close as possible to the original input.
Black-box Adversarial Attack Methods Query-basedScore-based Random Search SimBA[30], Parsimonious attack[31], Square attack[32], PPBA[33], BABIES[34] 111111111 111111111 Gradient Estimation NES[35], Bandits[36], ZO-signSGD[37], SignHunter[38], AdvFlow[39], NP-attack[40] 111111111 Combination 111111111 $\mathcal{CG}$-attack[41], MCG[42], PRGF[1], Subspace attack[43], BASES[44] 111111111 111111111 111111111 Decision-based Random Search Boundary attack[45], Evolutionary attack[46], GeoDA[47], SFA[48], Rays[49] 111111111 111111111 Gradient Estimation OPT[50], Sign-OPT[51], Triangle attack[52], HSJA[53], QEBA[54], NonLinear-BA[55], PSBA[56], CISA[2] Transfer-based| Data Perspective PGD[57], TI-FGSM[58], SI-FGSM[59], Admix[60], DI2-FGSM[61], SIA[3] 111111111 111111111 111111111 Optimization Perspective MI-FGSM[62], NI-FGSM[59], PI-FGSM[63], VT[64], RAP[65], LinBP[66], SGM[67], PGN[68] 111111111 Feature Perspective ILA[69], FIA[70], NAA[71] 111111111 111111111 111111111 111111111 111111111 111111111 Model Perspective 11(Tuning)11 RD[72], GhostNet[73], DRA[74], IAA[75], LGV[72], SWA[4], Bayesian attack[4] 111(Fusion)111 Logit ensemble[62], Loss ensemble[62], Longitudinal ensemble[73], CWA[76], AdaEA[77]